System of Equations using the Elimination Method
Looking for the problem worked out, not just the solution set. I am stuck but feel I am following the steps, just not sure where I am getting lost. Thank you.
9X + 2Y = 0
3X - 5Y = 17
I started by multiplying equation 1 by 5, and Equation 2 by 2 to make eliminate the Y's
5(9X+2Y)= 0 45X + 10Y = 0
51X = 34
2(3X-5Y)= 17 6X - 10Y = 34
51X/51 = 34/51 X= 0.666
9(.666 + 2Y) = 0
5.994 (rounded to 6) + 18Y = 0
6+18Y=0
-6 -6
18Y = -6
Y = .333
Perhaps the decimal is throwing me off?




Answer by ankor@dixie-net.com(22740)

I think you are on the right track, I will just show you the way I would do it.
9X + 2Y = 0
3X - 5Y = 17
I started by multiplying equation 1 by 5, and Equation 2 by 2 to make eliminate the Y's
5(9X+2Y)= 0 > 45X + 10Y = 0
:
2(3X-5Y)= 17 > 6X - 10Y = 34
:
taking these two equations
45x + 10y = 0
6x - 10y = 34
---------------Adding these equations eliminates y, find x
51x = 34
x = 34/51, reduces to 2/3
:
Find y using the 1st equation 9x + 2y = 0,replace x with 2/3
9(2/3) + 2y = 0
6 + 2y = 0
2y = -6
y = -6/2 reduces to -3
:
;
Check solutions in the 2nd equation; 3x - 5y = 17
3(2/3) - 5(-3) = 17
minus a minus is plus
2 + 15 = 17